Solve 8 sin? (w) – 10 sin(w) + 3 = 0 for all solutions ( <w < 27 W = Preview
Answer #1

solution ale have, W V U 8 sin? (W) - 10 sin(W) + 3 = 0 8 sen ?(0) - 6 sin (W)- 4 sin (W) + 3 =0 - 2 sin (W) (4 sin (W) - 3)